### v1

@@ Abstract
Graph-based Retrieval-augmented generation (RAG) has become a widely studied approach for improving the reasoning, accuracy, and factuality of Large Language Models. However, many existing graph-based RAG systems overlook the high cost associated with LLM token usage during graph construction, hindering large-scale adoption. To address this, we propose TERAG, a simple yet effective framework designed to build informative graphs at a significantly lower cost. Inspired by HippoRAG, we incorporate Personalized PageRank (PPR) during the retrieval phase, and we achieve at least 80% of the accuracy of widely used graph-based RAG methods while consuming only 3%-11% of the output tokens.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has emerged as an important framework to mitigate hallucinations and ground Large Language Models (LLMs) in external knowledge, enhancing their reliability in specialized domains like medicine [19], law [8], and education [10]. While traditional RAG retrieves unstructured text snippets, recent advancements have shifted towards graph-based RAG, which leverages knowledge graphs (KGs) to model structured relationships between information entities [23]. This structured approach enables more accurate multi-hop reasoning and provides greater transparency into the model’s decision-making process [5].
@@
To address this critical trade-off between cost and performance, we propose TERAG, a lightweight framework that minimizes LLM usage while leveraging their strengths in concept extraction. Instead of relying on multiple rounds of expensive LLM reasoning for graph construction, our method uses a few carefully designed prompts to extract multi-level concepts. These concepts are then structured into an effective knowledge graph using efficient, non-LLM methods, as illustrated in Figure 1 This process is illustrated with a real case from the 2WikiMultihopQA dataset. The two passages, while topically related, lack a direct connection that simple semantic retrieval could exploit. By extracting concepts from each passage and linking them with a Co_Occur edge, TERAG successfully connects them via key semantic information. This enables a successful retrieval for the question, “What is the death date of Lothair II’s mother? “—a query that would likely fail with retrieval methods based only on direct semantic similarity retrieval.
Figure 1: Graph structure of TERAG. The figure illustrates how lightweight concept extraction with LLMs, followed by non-LLM clustering and graph construction, leads to an efficient knowledge graph. In the graph, squares represent passages and circles represent concepts. Dark green circles denote directly extracted Named Entities, while dark blue circles denote extracted document-level concepts.
@@
  • We demonstrate that TERAG reduces ouput token usage by 89-97% compared to other widely-used graph-based RAG methods, offering a highly cost-effective solution.
@@ 2.0.1 Naive RAG
While pretrained language models have demonstrated a remarkable ability to internalize knowledge from training data, they possess significant limitations. Notably, their parametric memory is difficult to update, and they are prone to generating factually incorrect information, a phenomenon known as “hallucination“ [26, 20]. To mitigate these issues, Retrieval-Augmented Generation (RAG) was introduced to ground model responses in an external, non-parametric knowledge source [18]. Initial approaches, often termed “naive RAG,“ retrieve isolated text chunks based on vector similarity to a user’s query [9]. Although this method enhances factual accuracy, it fundamentally overlooks the interconnected nature of information. Consequently, it struggles with complex, multi-hop questions that require synthesizing insights from multiple, related documents [15]. This core limitation paved the way for more advanced methods like Graph-based RAG, which explicitly models the relationships between knowledge entities to enable more sophisticated reasoning.
@@ 2.0.2 Graph RAG
The emergence of Graph-based RAG addresses the primary limitation of its predecessors: the failure to model complex relationships between text chunks. At its core, Graph RAG leverages a pre-constructed knowledge graph, utilizing structural elements like nodes and paths to enrich the retrieval process and access more interconnected information [23]. This paradigm has inspired a wave of powerful systems, including Microsoft’s GraphRAG [5], HippoRAG [12], AutoSchemaKG [2] and ThinkOnGraph [27], many of which have demonstrated impressive performance and garnered significant industry attention, including Microsoft, NebulaGraph, and AntGroup [21, 22, 4].
The strength of these systems stems from a sophisticated and often costly indexing pipeline, where frameworks like GraphRAG [5], LightRAG [11], and MiniRAG [7] rely on Large Language Models (LLMs) to construct a knowledge graph from raw text. This process typically involves extracting entities and their connections as structured (subject, relation, object) triplets, and in some cases, generating node summaries or community reports [5]. While powerful, this deep integration of LLMs during indexing leads to substantial token consumption, creating a significant cost barrier for large-scale, real-world adoption.
@@ 3.2.3 Edge types
If two concept nodes $u,v\in V_{\mathrm{con}}$ appear in the same passage, we add bidirectional edges $(u\to v)$ and $(v\to u)$ labeled co_occur, avoiding duplicates to reduce graph density.
@@ 3.3 Objective
$\min T_{\text{total}}\quad\text{subject to}\quad\text{Accuracy}(RAG(G))\geq\delta,$
@@ 4.1.1 Named Entities
We adopt definition of Named Entity Recognition (NER) from Petasis et al. [25], where “a NE is a proper noun, serving as a name for something or someone,“ Specifically, our goal is to extract canonical mentions of salient entities. Formally, given a passage $p$, the NER model extracts a set of entity mentions
@@
All extracted entities across passages are aggregated into concept nodes.
$V_{\mathrm{con}}=\bigcup_{p\in V_{\mathrm{pas}}}\mathcal{E}(p).$
@@ 4.2 Graph Construction
If two nodes appear in the same passage, we create a bidirectional co_occur edge between them. This encodes local contextual associations between entities and concepts.
@@ 4.3.1 Query NER
Similar to entity and concept extraction in document processing, we use a few-shot prompt to extract named entities from the query. These extracted items are then matched against the concept node set $V_{\mathrm{con}}$ in the knowledge graph. We first attempt exact string matching; if no matches are found, we select the top-$3$ nodes with the highest semantic similarity (based on embeddings). The resulting matched nodes are used to construct a personalized dictionary that serves as the restart distribution for Personalized PageRank (PPR).
@@ 4.3.2 Personalized PageRank
We run a personalized PageRank algorithm on the knowledge graph, biased toward the query-relevant nodes identified in the previous step. Each matched node $u$ is assigned a weight based on two factors: semantic relevance and frequency. For nodes matched by exact string matching, the semantic weight is set to $1$. For nodes matched by semantic similarity, the semantic weight is given by the similarity score $s(u)$ between the query and the node embedding. In both cases, the frequency weight is defined as the inverse of the node frequency $f(u)$ in the corpus. Formally, the unnormalized weight $w(u)$ is defined as
$w(u)=\frac{s(u)}{f(u)},$
@@ 5.1.3 Baseline and Metrics
    $\text{EM}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}[\,y_{i}=y_{i}^{*}\,],$
@@
    $\text{Precision}_{i}$ $=\frac{|P_{i}\cap G_{i}|}{|P_{i}|},$
    $\text{Recall}_{i}$ $=\frac{|P_{i}\cap G_{i}|}{|G_{i}|}.$
@@
    $\text{F1}_{i}=\frac{2\cdot\text{Precision}_{i}\cdot\text{Recall}_{i}}{\text{Precision}_{i}+\text{Recall}_{i}},$
@@ 5.3 Token Consumption
AutoSchemaKG HotpotQA 5, 723, 7335,\,723,\,733 4, 915, 7964,\,915,\,796 10, 639, 52910,\,639,\,529
AutoSchemaKG 2WikiMultihopQA 3, 596, 6763,\,596,\,676 3, 176, 0953,\,176,\,095 6, 772, 7716,\,772,\,771
AutoSchemaKG MuSiQue 8, 960, 5028,\,960,\,502 7, 715, 9767,\,715,\,976 16, 676, 47816,\,676,\,478
LightRAG HotpotQA 38, 765, 23038,\,765,\,230 5, 862, 3635,\,862,\,363 44, 627, 59344,\,627,\,593
LightRAG 2WikiMultihopQA 34, 222, 64334,\,222,\,643 5, 288, 8065,\,288,\,806 39, 511, 44939,\,511,\,449
LightRAG MuSiQue 68, 500, 00068,\,500,\,000 18, 300, 00018,\,300,\,000 86, 800, 00086,\,800,\,000
MiniRAG HotpotQA 36, 909, 15036,\,909,\,150 14, 370, 41614,\,370,\,416 51, 279, 56651,\,279,\,566
MiniRAG 2WikiMultihopQA 13, 877, 88913,\,877,\,889 5, 906, 9845,\,906,\,984 19, 784, 87319,\,784,\,873
MiniRAG MuSiQue 62, 425, 40462,\,425,\,404 27, 552, 03127,\,552,\,031 89, 977, 43589,\,977,\,435
Note: Some data from the HippoRAG2 Paper [13]
@@
Table 4: Relative token consumption (TERAG = 100%) across datasets.
@@
MuSiQue $380.3$ 1, 160.81,\,160.8 $552.1$
LightRAG HotpotQA 1, 940.01,\,940.0 1, 041.61,\,041.6 1, 737.51,\,737.5
2WikiMultihopQA 2, 824.52,\,824.5 1, 434.41,\,434.4 2, 500.22,\,500.2
MuSiQue 2, 907.52,\,907.5 2, 753.12,\,753.1 2, 873.52,\,873.5
MiniRAG HotpotQA 1, 840.31,\,840.3 2, 553.32,\,553.3 1, 996.51,\,996.5
2WikiMultihopQA 1, 145.41,\,145.4 1, 602.11,\,602.1 1, 251.91,\,251.9
MuSiQue 2, 649.72,\,649.7 4, 145.04,\,145.0 2, 978.82,\,978.8
@@
Figure 2: Accuracy versus output token consumption on the 2WikiMultihopQA dataset. The upper-right region indicates higher accuracy with lower token consumption, representing more efficient performance.
@@ 6 Ablation Study
In this ablation study, we investigate the impact of our novel approach for calculating the Personalized PageRank restart vector. Our method combines semantic relevance and concept frequency, a key distinction from traditional PPR, which typically relies on a uniform distribution or inverse frequency alone. This analysis aims to quantify the effectiveness of our weighting scheme.
Table 5: Ablation study on PPR restart vector calculation. “New“ refers to our proposed method combining semantic relevance and frequency, while “Original“ uses only inverse frequency.
@@
We compared our modified PPR algorithm (TERAG + New) with a baseline that uses only the inverse of concept frequency as the restart distribution (TERAG + Original). As shown in Table 5, our approach consistently improves retrieval accuracy across all three datasets. Specifically, we observe an increase of 2-5 percentage points in both EM and F1 scores. This significant and stable improvement demonstrates that incorporating both semantic and frequency-based weighting is crucial for generating a more effective personalized dictionary, which in turn leads to a more accurate retrieval of relevant passages.
7 Conclusion
To address the high token consumption incurred by current Graph-RAG systems during knowledge graph construction, we propose the TERAG framework. This framework employs a concise process that constructs graph structures by directly extracting named entities and document-level concepts from text in a single pass, thereby avoiding the multiple and costly LLM reasoning calls common in existing methods.
Despite its lightweight design, TERAG achieves highly competitive retrieval accuracy on several standard multi-hop QA benchmarks, while reducing token overhead by one to two orders of magnitude. Our work demonstrates that by simplifying graph components and focusing only on passages, entities, and concepts, it is still possible to build a powerful Graph-RAG system. This indicates that a carefully designed lightweight graph construction pipeline can strike a better balance between efficiency and performance than approaches heavily reliant on LLMs.
For future work, we plan to explore extracting concepts at different levels of granularity and to incorporate additional engineering techniques to better identify the optimal balance between token consumption and model accuracy.

### v3

@@ Abstract
Graph-based Retrieval-augmented generation (RAG) has become a widely studied approach for improving the reasoning, accuracy, and factuality of Large Language Models (LLMs). However, many existing graph-based RAG systems overlook the high cost associated with LLM token usage during graph construction, hindering large-scale adoption. To address this, we propose TERAG, a simple yet effective framework designed to build informative graphs at a significantly lower cost. Inspired by HippoRAG, we incorporate Personalized PageRank (PPR) during the retrieval phase, and we achieve at least 80% of the accuracy of widely used graph-based RAG methods while consuming only 3%-11% of the output tokens. With its low token footprint and efficient construction pipeline, TERAG is well-suited for large-scale and cost-sensitive deployment scenarios.
@@ 1 Introduction
Retrieval-Augmented Generation has emerged as an important framework to mitigate hallucinations and ground Large Language Models in external knowledge, enhancing their reliability in specialized domains like medicine [18], law [8], and education [9]. While traditional RAG retrieves unstructured text snippets, recent advancements have shifted towards graph-based RAG, which leverages knowledge graphs (KGs) to model structured relationships between information entities [22]. This structured approach enables more accurate multi-hop reasoning and provides greater transparency into the model’s decision-making process [5].
@@
To address this critical trade-off between cost and performance, we propose TERAG, a lightweight framework that minimizes LLM usage while leveraging their strengths in concept extraction. The overall pipeline is shown in Figure 1. Instead of relying on multiple rounds of expensive LLM reasoning for graph construction, our method uses a few carefully designed prompts to extract multi-level concepts. These concepts are then structured into an effective knowledge graph using efficient, non-LLM methods, as illustrated in Figure 2 This process is illustrated with a real case from the 2WikiMultihopQA dataset. The two passages, while topically related, lack a direct connection that simple semantic retrieval could exploit. By extracting concepts from each passage and linking them with a co_occurrence edge, TERAG successfully connects them via key semantic information. This enables a successful retrieval for the question, “What is the death date of Lothair II’s mother? “—a query that would likely fail with retrieval methods based only on direct semantic similarity retrieval.
Figure 1: Overall pipeline of the proposed TERAG framework. The process consists of lightweight concept extraction with LLMs, followed by efficient non-LLM clustering and graph construction.
Figure 2: Graph structure of TERAG. The figure illustrates how lightweight concept extraction with LLMs, followed by non-LLM clustering and graph construction, leads to an efficient knowledge graph.
@@
  • We demonstrate that TERAG reduces output token usage by 89-97% compared to other widely-used graph-based RAG methods, offering a highly cost-effective solution.
@@ 2.0.1 Naive RAG
While pretrained language models have demonstrated a remarkable ability to internalize knowledge from training data, they possess significant limitations. Notably, their parametric memory is difficult to update, and they are prone to generating factually incorrect information, a phenomenon known as ”hallucination” [25, 19]. To mitigate these issues, Retrieval-Augmented Generation was introduced to ground model responses in an external, non-parametric knowledge source [17]. Initial approaches, often termed “naive RAG,“ retrieve isolated text chunks based on vector similarity to a user’s query. Although this method enhances factual accuracy, it fundamentally overlooks the interconnected nature of information. Consequently, it struggles with complex, multi-hop questions that require synthesizing insights from multiple, related documents [14]. This core limitation paved the way for more advanced methods like Graph-based RAG, which explicitly models the relationships between knowledge entities to enable more sophisticated reasoning.
@@ 2.0.2 Graph RAG
The emergence of Graph-based RAG addresses the primary limitation of its predecessors: the failure to model complex relationships between text chunks. Graph RAG leverages a pre-constructed knowledge graph, utilizing structural elements like nodes and paths to enrich the retrieval process and access more interconnected information [22]. This paradigm has inspired a wave of powerful systems, including Microsoft’s GraphRAG [5], HippoRAG [11], AutoSchemaKG [2] and ThinkOnGraph [26], many of which have demonstrated impressive performance and garnered significant industry attention, including Microsoft, NebulaGraph, and AntGroup [20, 21, 4].
The strength of these systems stems from a sophisticated and often costly indexing pipeline, where frameworks like GraphRAG [5], LightRAG [10], and MiniRAG [7] rely on Large Language Models (LLMs) to construct a knowledge graph from raw text. This process typically involves extracting entities and their connections as structured (subject, relation, object) triplets, and in some cases, generating node summaries or community reports [5]. While powerful, this deep integration of LLMs during indexing lead to substantial token consumption, creating a significant cost barrier for large-scale, real-world adoption.
@@ 3.2.3 Edge types
If two concept nodes $u,v\in V_{\mathrm{con}}$ appear in the same passage, we add bidirectional edges $(u\to v)$ and $(v\to u)$ labeled co_occurrence, avoiding duplicates to reduce graph density.
@@ 3.3 Objective
$\min T_{\text{total}}\quad\text{subject to}\quad\text{Accuracy}(RAG(G))\geq\delta$ (1)
@@ 4.1.1 Named Entities
We adopt the definition of Named Entity Recognition (NER) following Petasis et al. [24], where “a Named Entity (NE) is a proper noun serving as a name for something or someone.” Our goal is to extract canonical mentions of salient entities.
Formally, given a passage $p$, the NER model extracts a set of entity mentions:
@@
All extracted entities across passages are aggregated into a unified concept node set:
$V_{\mathrm{ent}}=\bigcup_{p\in V_{\mathrm{pas}}}\mathcal{E}(p),$
where $V_{\mathrm{pas}}$ denotes the set of passage nodes in the knowledge graph.
@@ 4.2 Graph Construction
If two nodes appear in the same passage, we create a bidirectional co_occurrence edge between them. This encodes local contextual associations between entities and concepts.
@@ 4.3.1 Query NER
Similar to entity and concept extraction in document processing, we use a few-shot prompt to extract named entities from the query. These extracted items are then matched against the concept node set $V_{\mathrm{con}}$ in the knowledge graph. We first attempt exact string matching; if no matches are found, we select the top-$3$ nodes with the highest semantic similarity (based on embeddings). The resulting matched nodes are used to construct a personalized dictionary that serves as the restart distribution for PPR.
@@ 4.3.2 Personalized PageRank
We run a PPR algorithm on the knowledge graph, biased toward the query-relevant nodes identified in the previous step. Each matched node $u$ is assigned a weight based on two factors: semantic relevance and frequency. For nodes matched by exact string matching, the semantic weight is set to $1$. For nodes matched by semantic similarity, the semantic weight is given by the similarity score $s(u)$ between the query and the node embedding. In both cases, the frequency weight is defined as the inverse of the node frequency $f(u)$ in the corpus. Formally, the unnormalized weight $w(u)$ is defined as
$w(u)=\frac{s(u)}{f(u)},$ (3)
@@ 5.1.3 Baseline and Metrics
    $\text{EM}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}[\,y_{i}=y_{i}^{*}\,]$ (6)
@@
    $\text{Precision}_{i}$ $=\frac{|P_{i}\cap G_{i}|}{|P_{i}|},$ (7)
    $\text{Recall}_{i}$ $=\frac{|P_{i}\cap G_{i}|}{|G_{i}|}.$ (8)
@@
    $\text{F1}_{i}=\frac{2\cdot\text{Precision}_{i}\cdot\text{Recall}_{i}}{\text{Precision}_{i}+\text{Recall}_{i}}$ (9)
@@ 5.3 Token Consumption
AutoSchemaKG HotpotQA $5,723,733$ $4,915,796$ $10,639,529$
AutoSchemaKG 2WikiMultihopQA $3,596,676$ $3,176,095$ $6,772,771$
AutoSchemaKG MuSiQue $8,960,502$ $7,715,976$ $16,676,478$
LightRAG HotpotQA $38,765,230$ $5,862,363$ $44,627,593$
LightRAG 2WikiMultihopQA $34,222,643$ $5,288,806$ $39,511,449$
LightRAG MuSiQue $68,500,000$ $18,300,000$ $86,800,000$
MiniRAG HotpotQA $36,909,150$ $14,370,416$ $51,279,566$
MiniRAG 2WikiMultihopQA $13,877,889$ $5,906,984$ $19,784,873$
MiniRAG MuSiQue $62,425,404$ $27,552,031$ $89,977,435$
Note: Some data from the HippoRAG2 paper [12].
@@
Table 4: Relative token consumption across datasets. Manually Set TERAG = 100%.
@@
MuSiQue $380.3$ $1,160.8$ $552.1$
LightRAG HotpotQA $1,940.0$ $1,041.6$ $1,737.5$
2WikiMultihopQA $2,824.5$ $1,434.4$ $2,500.2$
MuSiQue $2,907.5$ $2,753.1$ $2,873.5$
MiniRAG HotpotQA $1,840.3$ $2,553.3$ $1,996.5$
2WikiMultihopQA $1,145.4$ $1,602.1$ $1,251.9$
MuSiQue $2,649.7$ $4,145.0$ $2,978.8$
@@
Figure 3: Accuracy versus output token consumption on the 2Wiki dataset. The upper-right region indicates higher accuracy with lower token consumption.
@@ 6 Ablation Study
Effect of Restart Vector Design.
We examine the impact of our new PPR restart vector formulation, which integrates both semantic relevance and concept frequency. Unlike traditional PPR methods that rely solely on uniform or inverse-frequency distributions, our approach aims to produce a more informative restart vector that better captures contextual importance.
Table 5: Ablation study on PPR restart vector calculation. “New“ denotes our method combining semantic relevance and frequency, while “Original“ uses only inverse frequency.
@@
As shown in Table 5, our restart vector design consistently improves retrieval accuracy across all three datasets, yielding gains of 2–5 percentage points in both EM and F1. These results highlight the importance of combining semantic and frequency information to construct a more discriminative restart distribution, leading to more precise passage retrieval.
Model Variation Analysis.
We also assessed the influence of switching the answer generation model while keeping the same retrieval graph. Replacing the default model with LLaMA-3.2 11B caused a slight performance drop, while Qwen2 235B achieved accuracy similar to the LLaMA-3 70B but introduced noticeably higher latency. These observations suggest that the dominant performance factor in our framework is the model’s NER accuracy; improvements in generic reasoning or generation capabilities do not yield proportional gains in retrieval quality.
7 Conclusion and Future Work
To address the high token consumption incurred by current Graph-RAG systems during knowledge graph construction, we propose the TERAG framework. TERAG constructs graph structures by directly extracting named entities and document-level concepts from text in a single pass, thereby eliminating the multiple and costly LLM reasoning calls required by prior approaches.
Despite its lightweight design, TERAG achieves highly competitive retrieval accuracy on multiple multi-hop QA benchmarks while reducing token overhead by one to two orders of magnitude. In addition, because concept extraction, concept filtering, and graph linking are largely independent, these components can be parallelized efficiently. As a result, TERAG not only minimizes indexing-related token cost but also has the potential to significantly shorten graph construction latency in large-scale deployments. Overall, our findings demonstrate that a carefully designed lightweight graph construction pipeline can achieve a more favorable efficiency–performance trade-off than graph-RAG systems that heavily rely on LLMs.
For future work, we plan to explore extracting concepts with controllable levels of granularity, enabling multi-layered knowledge graphs that better capture semantic structure and improve retrieval accuracy. We also aim to incorporate additional engineering strategies to further optimize the balance between token efficiency and model effectiveness.
Finally, due to TERAG’s simple and adaptable pipeline, we expect its performance to remain stable under evolving data streams or more heterogeneous corpora. However, the lack of widely recognized benchmarks for such dynamic and heterogeneous retrieval scenarios limits empirical evaluation. Future research will involve designing appropriate datasets and evaluation protocols to more systematically validate these capabilities.